### v1

# Complexities of Theories of Profinite Subgroups of $S_{\omega}$ via Tree Presentations

Abstract
Although $S_{\omega}$ (the group of all permutations of $\mathbb{N}$) is size continuum, both it and its closed subgroups can be presented as the set of paths through a countable tree. The subgroups of $S_{\omega}$ that can be presented this way with finite branching trees are exactly the profinite ones. We use these tree presentations to find upper bounds on the complexity of the existential theories of profinite subgroups of $S_{\omega}$, as well as to prove sharpness for these bounds. These complexity results enable us to distinguish a simple subclass of profinite groups, those with orbit independence, for which we find an upper bound on the complexity of the entire first order theory.
Keywords:
Computable structure theory Permutation groups Profinite groups Tree presentations.
1 Introduction
Traditional computable structure theory deals only with countable structures. As a result, it cannot be used to study most subgroups of $S_{\omega}$ (the group of all permutations of $\mathbb{N}$). However, as described in Section 2, a large class of these subgroups (specifically the closed subgroups) can be presented as the set of paths through a countable tree. We will focus our attention on the subgroups that can be presented as the paths through a finite branching tree (the compact subgroups), which are exactly the profinite ones. In [5] the author uses such a presentation to study the absolute Galois group of $\mathbb{Q}$, which is indeed a profinite group that can be viewed as a subgroup of $S_{\omega}$ after fixing an enumeration of the algebraic closure of $\mathbb{Q}$.
Much interest in profinite groups stems from their connection to Galois theory. As shown in [8], every profinite group is the Galois group of some field extension. However, the purpose of this paper is to examine profinite groups simply as groups in their own right. It can be difficult to get an effective handle on uncountable groups, but when such a group acts on $\mathbb{N}$ by permutations we are given the opportunity to do so. Thus, we restrict our attention to the profinite subgroups of $S_{\omega}$.
Effective notions for profinite groups within the context of Galois theory were examined in [4] and further in [3]. Following this work, effective notions for profinite groups in general were studied in [7]. The authors of [3] and [7] define a profinite group to be recursively profinite if it is isomorphic to the inverse limit of a uniformly computable sequence of finite groups and surjective homomorphisms. As we will see in Proposition 4, a profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\omega}$ such that $T_{G}$ (the tree that represents $G$) is computable. More recently, effectively closed subgroups of $S_{\omega}$ were examined in [2].
In Section 3, we use tree presentations to determine bounds on the complexity of the existential theory of profinite subgroups of $S_{\omega}$. Note that since all elements of $S_{\omega}$ are functions from $\mathbb{N}\to\mathbb{N}$, an existential sentence about a subgroup is a $\Sigma^{1}_{1}$ statement. However, we will see that the existential theory of any profinite subgroup $G$ of $S_{\omega}$ is $\Sigma^{0}_{2}$ relative to the degree of $T_{G}$. Additionally, if $G$ has orbit independence, then the existential theory is $\Sigma^{0}_{1}$ relative to the degree of $T_{G}$. We will also show that these bounds are sharp. Specifically, there exists a profinite $G$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete, and such a $G$ without orbit independence such that the existential theory is $\Sigma^{0}_{2}$ complete. Last, we show that the (entire) first order theory of a profinite $G$ with orbit independence is $\Delta^{0}_{2}$ relative to the degree of $T_{G}$.
2 Tree Presentations
Definition 1
Let $G$ be a subgroup of $S_{\omega}$. We define the tree $T_{G}$ to be the subtree of $\mathbb{N}^{<\omega}$ containing all initial segments of elements of $G$. That is,
$T_{G}:=\{\tau\in\mathbb{N}^{<\omega}:(\exists g\in G,n\in\mathbb{N})[\tau=g(0)g(1)\cdots g(n)]\}$
where $m\in\mathbb{N}$ is mapped to $g(m)$ under $g$. We define the ordering of $T_{G}$ via initial segments and write $\tau\sqsubset\sigma$ if $\tau$ is an initial segment of $\sigma$.
It should be noted that $T_{G}$ will have no terminal nodes. That is, every element of $T_{G}$ is an initial segment of another.
Definition 2
Let $G$ be a subgroup of $S_{\omega}$. We define the degree of $T_{G}$ to be the join of the Turing degrees of
[…]
  • A branching function $Br:T_{G}\to\mathbb{N}\cup\{\infty\}$ such that $Br(\tau)$ is equal to the number of direct successors of $\tau$ in $T_{G}$.
We denote the degree of $T_{G}$ as $\deg(T_{G})$.
We will focus on groups where $T_{G}$ is finite branching, in which case the range of $Br$ will be a subset of $\mathbb{N}$. It should be noted that $\deg(T_{G})$ is not invariant under group isomorphism in that it is possible to have $G\cong G^{\prime}$ with $\deg(T_{G})\neq\deg(T_{G^{\prime}})$.
Definition 3
Given a tree $T\subset\mathbb{N}^{<\omega}$, we define $[T]$ to be the set of all paths through $T$. We endow $[T]$ with the standard product topology in which the basic clopen sets are those of the form $\{f\in\mathbb{N}^{\omega}:\tau\sqsubset f\}$ for some $\tau\in T$.
It is clear that every element of $G$ is represented as a path through $T_{G}$. In particular, the function $i:G\to[T_{G}]$ defined by
$i(g)=g(0)g(1)\cdots$
[…]
Proposition 1
Let $G$ be a subgroup of $S_{\omega}$. The map $i:G\to[T_{G}]$ is a bijection if and only if $i(G)$ is a closed subset of $[T_{G}]$. ∎
We say that $G$ is a closed group when $i(G)$ is closed. Thus, the subgroups of $S_{\omega}$ that can be represented as the paths through this type of tree are exactly the closed ones. Additionally, we say that $G$ is a compact group if $i(G)$ is compact.
Definition 4
A topological group is called profinite if it is isomorphic to the inverse limit of an inverse system of discrete finite groups.
The following proposition yields a simple topological definition for profinite groups.
Proposition 2 (Folklore; see e.g. Theorem 3.7 from [6])
A topological group is profinite if and only if it is Hausdorff, compact, and totally disconnected.∎
Definition 5
Given a subgroup $G$ of $S_{\omega}$ and $n\in\mathbb{N}$, we define the orbit of $n$ under $G$ as
$\text{orb}_{G}(n):=\{g(n)\in\mathbb{N}:g\in G\}.$
The following proposition is also folklore, but we give a brief proof.
Proposition 3
Let $G$ be a subgroup of $S_{\omega}$. The following are equivalent:
  1. (1)
    $G$ is compact,
  2. (2)
    $G$ is closed and all orbits under $G$ are finite,
  3. (3)
    $G$ is profinite.
Proof
Suppose that $G$ is compact . Since our topology is Hausdorff, we have that $G$ is closed. By Proposition 1, we have that $i(G)=[T_{G}]$. Assume towards a contradiction that there is some $n\in\mathbb{N}$ with $\text{orb}_{G}(n)$ infinite. Let $\{\tau_{i}\}_{i\in\mathbb{N}}$ be the (infinite) collection of all elements of $T_{G}$ of length $n+1$. Note that $\{\{f:\tau_{i}\sqsubset f\}\}_{i\in\mathbb{N}}$ is an open cover of $[T_{G}]=i(G)$ with no finite subcover, which contradicts that $G$ is compact. Hence, $(1)\implies(2)$.
If $G$ is closed and all orbits in $G$ are finite, then it follows that $G$ is compact as a consequence of König’s lemma. Hence, $(2)\implies(1)$.
The topology we have defined is Hausdorff and totally disconnected. If $G$ is compact, then $G$ is also profinite by Proposition 2. Hence, $(3)\iff(1)$. ∎
We have that all profinite subgroups of $S_{\omega}$ will have countably many orbits (all of which are finite). We fix an enumeration of these orbits as follows:
Definition 6
Let $G$ be a profinite subgroup of $S_{\omega}$. Define $\{O_{G,i}\}_{i\in\mathbb{N}}$ so that $O_{G,0}=\text{orb}_{G}(0)$ and $O_{G,n+1}$ is the orbit of the least natural number not in any $O_{G,m}$ with $m\leq n$.
We can use these orbits to define finite approximations of $G$ up to the first $k\in\mathbb{N}$ many orbits.
Definition 7
Let $G$ be a profinite subgroup of $S_{\omega}$. Given $g\in G$ and $k\in\mathbb{N}$, define $g_{k}=g\upharpoonright\bigcup_{i\leq k}O_{G,i}$. Define
$G_{k}:=\{g_{k}:g\in G\}.$
We can also define the restriction of $G$ to only the $k$th orbit.
Definition 8
Let $G$ be a profinite subgroup of $S_{\omega}$. Define
$H_{k}:=\{g\upharpoonright O_{G,k}:g\in G\}.$
Note that both $G_{k}$ and $H_{k}$ are finite groups for all $k\in\mathbb{N}$. Additionally, they are both uniformly computable given $T_{G}$.
Definition 9
Let $G$ be a profinite subgroup of $S_{\omega}$. We say that $G$ has orbit independence if it is isomorphic to the Cartesian product of all $H_{k}$. That is,
$G\cong\prod_{k\in\mathbb{N}}H_{k}.$
For an example of a profinite subgroup that does not have orbit independence, consider
[…]
where $1_{G}$ denotes the identity permutation. Note that $H_{0}\cong H_{1}\cong C_{2}$ (the cyclic group on $2$ elements) and $H_{n}$ is trivial for all $n>1$. Thus, $\prod H_{k}\cong C_{2}\times C_{2}$ but $G\cong C_{2}$. As we shall see, groups with orbit independence tend to be simpler to work with.
It should be noted that our definition for the degree of $T_{G}$ is compatible with the notion of a recursively profinite group used by La Roche and Smith.
Definition 10 ([3, 7])
A profinite group $P$ is called recursively profinite if there exists a uniformly computable sequence $\{P_{n},\pi_{n}\}_{n\in\mathbb{N}}$ such that each $P_{n}$ is a finite group, each $\pi_{n}$ is a surjective homomorphism from $P_{n+1}$ to $P_{n}$, and $P$ is isomorphic to the inverse limit of the sequence.
Proposition 4
A profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\omega}$ with $T_{G}$ computable.
Proof
Suppose that $P\cong G$ with $T_{G}$ computable. Defining $\nu_{n}:G_{n+1}\to G_{n}$ (with $G_{n}$ as in definition 8) so that $\nu_{n}(g_{n+1})=g_{n}$, we get that $\{G_{n},\nu_{n}\}_{n\in\mathbb{N}}$ is a uniformly computable sequence as required in definition 10 whose inverse limit is isomorphic to $P$.
For the other direction, suppose that $\{P_{n},\pi_{n}\}_{n\in\mathbb{N}}$ is as in definition 10. For each $n\in\mathbb{N}$, let $N_{n}$ be a natural number such that $P_{n}$ is isomorphic to a subgroup of $S_{N_{n}}$ (the group of permutations of $\{0,...,N_{n}-1\}$). Define $f_{0}:P_{0}\to S_{N_{0}}$ such that $f_{0}$ is a group embedding. Given $f_{n}$, define $f_{n+1}:P_{n+1}\to S_{N_{n+1}}$ such that $f_{n+1}$ is a group embedding and “respects” $\pi_{n+1}$ in the sense that if $\pi_{n+1}(p_{n+1})=p_{n}$, then $f_{n+1}(p_{n+1})\upharpoonright N_{n}=f_{n}(p_{n})$. Define $G$ to be the set of $g\in S_{\omega}$ such that for all $n\in\mathbb{N}$, there exists a $p\in P_{n}$ with $f_{n}(p)=g\upharpoonright N_{n}$. We have that $G\cong P$ and that $T_{G}$ is exactly the set
$\left\{\tau\in\mathbb{N}^{<\omega}:\left(\exists n\in\mathbb{N},p\in P_{n},m<N_{n}\right)[\tau=f_{n}(p)(0)f_{n}(p)(1)\cdots f_{n}(p)(m)]\right\}.$
It is clear that the domain and branching function of $T_{G}$ are computable, hence we have that $T_{G}$ is computable. ∎
3 Complexity of Theories
We now consider the complexity of the existential theory of a profinite subgroup of $S_{\omega}$. To do so, we must first establish a few lemmas.
Definition 11
A positive formula is a first order formula that can be expressed without the use of any negation symbols. A negative formula is the negation of a positive formula.
Lemma 1
Let $G$ be a profinite subgroup of $S_{\omega}$ and let $\alpha^{+}$ be a positive formula in the language of groups. If $k<l$, then
$G_{l}\models\alpha^{+}(\bar{g}_{l})\implies G_{k}\models\alpha^{+}(\bar{g}_{k})$
for any $\bar{g}\in G^{<\omega}$.
Proof
If $\alpha^{+}$ is quantifier free, than it is expressible as a disjunction of conjunctions of atomic formulas. Recall that an atomic formula in the language of groups is equivalent to the statement that some word is equal to the identity. If $G_{l}\models\alpha^{+}(\bar{g}_{l})$, then one of the disjuncts must hold which just means that some collection of words $W_{1}(\bar{g}_{l}),...,W_{n}(\bar{g}_{l})$ over the alphabet $\{x,x^{-1}:x\in\bar{g}_{l}\}$ are all equal to $1_{G_{l}}$. Since each element of $\bar{g}_{k}$ is an initial segment of an element of $\bar{g}_{l}$, we must also have that $W_{1}(\bar{g}_{k})=\cdots=W_{n}(\bar{g}_{k})=1_{G_{k}}$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$.
Now suppose that the result holds for all $\Sigma_{n}$ and $\Pi_{n}$ formulas. If $\alpha^{+}$ is $\Sigma_{n+1}$, then we have $\alpha^{+}\equiv\exists\bar{x}\beta^{+}$ where $\beta^{+}$ is a $\Pi_{n}$ positive formula. If $G_{l}\models(\exists\bar{x})\beta^{+}(\bar{g}_{l},\bar{x})$ then there exists some $\bar{h}\in G^{<\omega}$ such that $G_{l}\models\beta^{+}(\bar{g}_{l},\bar{h}_{l}).$ Since the result holds of $\Pi_{n}$ formulas, we get that $G_{k}\models\beta^{+}(\bar{g}_{k},\bar{h}_{k})$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$.
If $\alpha^{+}$ is $\Pi_{n+1}$, then we have $\alpha^{+}\equiv\forall\bar{x}\beta^{+}$ where $\beta^{+}$ is $\Sigma_{n}$. If $G_{l}\models(\forall\bar{x})\beta^{+}(\bar{g}_{l},\bar{x})$ then $G_{l}\models\beta^{+}(\bar{g}_{l},\bar{h}_{l})$ for all $\bar{h}\in G^{<\omega}$. Again, since the result holds for all $\Sigma_{n}$ formulas we have that $G_{k}\models\beta^{+}(\bar{g}_{k},\bar{h}_{k})$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$. ∎
Corollary 1
Let $G,k$ and $l$ be as in the previous Lemma. If $\alpha^{-}$ is a negative formula, then
$G_{k}\models\alpha^{-}(\bar{g}_{k})\implies G_{l}\models\alpha^{-}(\bar{g}_{l})$
for any $\bar{g}\in G^{<\omega}$. ∎
Lemma 2
Let $G$ be a profinite subgroup of $S_{\omega}$. If $\alpha$ is quantifier free, then $G\models\alpha(\bar{g})$ if and only if $G_{k}\models\alpha(\bar{g}_{k})$ for all sufficiently large $k\in\mathbb{N}$.
Proof
For the base case, let $\alpha$ be atomic. We have that $\alpha(\bar{x})\equiv W(\bar{x})=1$ where $W(\bar{x})$ is a word over $\{x,x^{-1}:x\in\bar{x}\}$ and $1$ is the group identity symbol. Clearly, if $W(\bar{g})=1_{G}$ then $W(\bar{g}_{k})=1_{G_{k}}$ for all $k$. On the other hand, if $W(\bar{g}_{k})=1_{G_{k}}$ for sufficiently large $k$ then we have by the previous lemma that $W(\bar{g}_{k})=1_{G_{k}}$ for all $k\in\mathbb{N}$. Thus given any $n\in\mathbb{N}$ and a large enough $l$ such that $n\in\text{dom}(\bar{g}_{l})$, we have that $W(\bar{g}_{l})$ maps $n$ to $n$. Hence, $W(\bar{g})=1_{G}$.
Negative Step:
Let $\alpha\equiv\neg\beta$ with $\beta$ atomic. Suppose $G\models\neg\beta(\bar{g})$. This gives that there is a $k$ such that $G_{k}\models\neg\beta(\bar{g}_{k})$. By the previous lemma, we must have that $G_{l}\models\neg\beta(\bar{g}_{l})$ for all $l\geq k$ and thus for all sufficiently large $l$.
Now suppose that $G_{k}\models\neg\beta(\bar{g}_{k})$ for sufficiently large $k$. There must only be finitely many $k$ such that $G_{k}\models\beta(\bar{g}_{k})$. Thus, from the base case, we have that $G\models\neg\beta(\bar{g})$.
Conjunctive/Disjunctive Step:
If the statement holds for $\beta_{1}$ and $\beta_{2}$, then it is clear that it holds for $\beta_{1}\,\&\,\beta_{2}$ as well. If the statement holds for $\beta_{1}$ or for $\beta_{2}$, then it is clear that it holds for $\beta_{1}\vee\beta_{2}$ as well. ∎
Lemma 3
Every atomic sentence in the language of groups is true in every group.
Proof
Every such sentence has the form
$1^{n}=1^{m}$
with $n,m\in\mathbb{Z}$.∎
The following lemma gives an example of the power of orbit independence.
Lemma 4
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. Let $\alpha$ be an existential sentence in the language of groups. We have that $G\models\alpha$ if and only if $G_{k}\models\alpha$ for some $k\in\mathbb{N}$.
Proof
We have that $\alpha\equiv\exists\bar{x}\beta$ where $\beta$ is a quantifier free formula. If $G\models\alpha$ then there is some $\bar{g}\in G^{<\omega}$ such that $G\models\beta(\bar{g})$. Thus, Lemma 2 gives that there is a $k$ with $G_{k}\models\beta(\bar{g}_{k})$ and so $G_{k}\models\alpha$.
Now suppose that $G_{k}\models\alpha$. We have that for some $\bar{\gamma}\in G_{k}^{<\omega}$, $G_{k}\models\beta(\bar{\gamma})$. Define an embedding $f:G_{k}\to G$ such that $f(\gamma)\upharpoonright\text{dom}(\gamma)=\gamma$ and $f(\gamma)$ is just the identity on all orbits $O_{G,l}$ with $l>k$. Note that since $G$ has orbit independence, we will in fact have that $f(\gamma)\in G$.
Note that
$\beta(\bar{x})\equiv\bigvee_{i}\bigwedge_{j}\beta^{+}_{i,j}(\bar{x})\,\&\,\beta^{-}_{i,j}(\bar{x})$
where each $\beta^{+}$ is atomic and each $\beta^{-}$ negated atomic. Since $G_{k}\models\beta(\bar{\gamma})$, there is some $i$ such that $G_{k}\models\bigwedge_{j}\beta^{+}_{i,j}(\bar{\gamma})\,\&\,\beta^{-}_{i,j}(\bar{\gamma}).$ It is clear that any negated atomic formula that holds of $\bar{\gamma}$ in $G_{k}$ will also hold of $f(\bar{\gamma})$ in $G$. Thus if $G\nvDash\alpha$, there would have to be some $j$ such that $G\nvDash\beta^{+}_{i,j}(f(\bar{\gamma}))$. However since we are assuming $G_{k}\models\beta^{+}_{i,j}(\bar{\gamma})$, there would have to be some $l>k$ such that $H_{l}\models\neg\beta^{+}_{i,j}(f(\bar{\gamma})\upharpoonright O_{G,l})$. Since $\bar{\gamma}$ is just the identity on all orbits $O_{G,l}$ with $l>k$, we would have $H_{l}\models\neg\beta^{+}_{i,j}(\bar{1})$. The formula $\beta^{+}_{i,j}(\bar{1})$ is an atomic sentence in the language of groups, and so by Lemma 3 we get that $H_{l}\models\beta^{+}_{i,j}(\bar{1})$. Hence, we must have that $G\models\alpha$. ∎
Theorem 3.1
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. The existential theory of $G$ is $\Sigma^{0}_{1}$ relative to $\deg(T_{G})$.
Proof
Let $\alpha$ be an existential sentence. By the previous lemma, $G\models\alpha$ if and only if
$(\exists k)[G_{k}\models\alpha]$
which is $\Sigma^{0}_{1}$ relative to $\deg(T_{G})$. ∎
The following proposition gives that the above theorem is sharp.
Proposition 5
There exists a profinite subgroup $G$ of $S_{\omega}$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete.
Proof
By the previous theorem, we need only build a $G$ with $T_{G}$ computable such that the existential theory codes $\emptyset^{\prime}$. Define the formula $\alpha_{n}$ for all $n\in\mathbb{N}$ by
$\alpha_{n}:=(\exists x)[x\neq 1\,\&\,x^{p_{n}}=1]$
where $1$ is the identity element and $\{p_{n}\}_{n\in\mathbb{N}}$ is the sequence of all primes. We build $G$ such that $G\models\alpha_{n}$ if and only if $n\in\emptyset^{\prime}$.
Construction
Stage 00:
Define $O_{G,0}=\{0\}$ and define $H_{0}$ to be the trivial group.
Stage $s$:
Let $N_{s}\in\mathbb{N}$ be the least not in any $O_{G,i}$ with $i<s$. Find the least $e\leq s$ such that $\Phi_{e,s}(e)\downarrow$ and $G_{s-1}\models\neg\alpha_{e}$. If no such $e$ exists, define $O_{G,s}=\{N_{s}\}$ and $H_{s}$ to be the trivial group. If there is such an $e$, define $O_{G,s}=\{N_{s},N_{s}+1,...,N_{s}+p_{e}-1\}$ and define $H_{s}$ to be cyclic on $O_{G,s}$.
Verification
Since each $G_{s}$ is computable, it is clear that the tree $T_{G}$ is computable. If $n\notin\emptyset^{\prime}$, then no $H_{s}$ will be of size $p_{n}$. Thus, no element has order $p_{n}$, which gives $G\models\neg\alpha_{n}$. If $n\in\emptyset^{\prime}$, then there will come a stage $t$ in which $n$ is the least such that $\Phi_{n,t}(e)\downarrow$ and there is currently no $H_{s}$ of size $p_{n}$. We will then make $H_{t}$ cyclic and of size $p_{n}$ which will assure that $G\models\alpha_{n}$. ∎
Theorem 3.2
Let $G$ be any profinite subgroup of $S_{\omega}$ (not necessarily with orbit independence). The existential theory of $G$ is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$.
Proof
Suppose $\alpha=\exists\bar{x}\beta$ with $\beta$ quantifier free. By Lemma 2, given $\bar{g}\in G^{<\omega}$ we have that $G\models\beta(\bar{g})$ if and only if $G_{k}\models\beta(\bar{g}_{k})$ for all but finitely many $k\in\mathbb{N}$. Let $T_{\beta}$ be the subset of $T_{G}$ defined by
$T_{\beta}:=\{\tau\in T_{G}:G_{l(\tau)}\models\beta(\tau))\}$
where $l(\tau)$ is defined as the natural number such that $\tau\in G_{l(\tau)}$. Note that $T_{\beta}$ is computable given $T_{G}$. We have that $G\models\alpha$ if and only if
$(\exists\bar{\tau}\in T_{G}^{<\omega})(\forall k\geq l(\bar{\tau}))\left[\bigvee_{\bar{\sigma}\in G_{k}^{<\omega}}\left(\bar{\tau}\sqsubseteq\bar{\sigma}\,\&\,\bigwedge_{\bar{\tau}\sqsubseteq\bar{\rho}\sqsubseteq\bar{\sigma}}\bar{\rho}\in T_{\beta}\right)\right]$
which is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$ (recall that $\deg(T_{G})$ computes the branching function $T_{G}$, and can thus compute the elements of each $G_{k}$). ∎
The following proposition gives that the above theorem is sharp.
Proposition 6
There exists a profinite subgroup $G$ of $S_{\omega}$ (without orbit independence) with $T_{G}$ computable such that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
Proof
Recall that the set $\text{Fin}=\{e\in\mathbb{N}:|W_{e}|<\infty\}$ (where $W_{e}$ is the domain of $\Phi_{e}$, the $e$th Turing program) is $\Sigma^{0}_{2}$ complete. Let $\{p_{n}\}_{n\in\mathbb{N}}$ be the sequence of all primes. Given $n\in\mathbb{N}$, define the formula
$\alpha_{n}:=(\exists x)[x\neq 1\,\&\,x^{p_{n}}=1].$
We construct $G$ so that
$G\models\alpha_{n}\iff n\in\text{Fin}.$
This, along with Lemma 3.2, ensures that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
We construct $G$ in stages, defining $G_{s}$ at stage $s$ of the construction. At stages of the form $s=\langle n,m\rangle$, we work toward making sure that $G$ will model $\alpha_{n}$ just if $W_{n}$ is finite. Specifically, if $|W_{n,m+1}|>|W_{n,m}|$ then we make sure that if $g\in G_{s-1}$ with $g\neq 1_{G_{s-1}}$, then any $g^{\prime}\in G_{s}$ with $g\sqsubset g^{\prime}$ has $g^{\prime p_{n}}\neq 1_{G_{s}}$. We also create a new element not equal to $1_{G_{s}}$ that is of order $p_{n}$. If $W_{n,m+1}=W_{n,m}$, then we define $H_{s}$ to be the trivial group (thus if $g\sqsubset g^{\prime}$ with $g\in G_{s-1}$ and $g^{\prime}\in G_{s}$, then $g^{\prime p_{n}}=1_{G_{s}}$ if and only if $g^{p_{n}}=1_{G_{s-1}}$).
Construction
Define a bijection $\langle\rangle:\mathbb{N}^{2}\to\mathbb{N}$ such that $0=\langle 0,0\rangle$ and $\langle n,m\rangle<\langle n,m+1\rangle$ for all $n,m\in\mathbb{N}$. Define $l_{0}=0$. For all $s>0$, define $l_{s}$ to be the least natural number not in $O_{G,s-1}$.
Stage $0=\langle 0,0\rangle$:
[…]
Stage $s=\langle n,0\rangle$ with $n>0$:
Define $O_{G,s}=\{l_{s},l_{s}+1,...,l_{s}+p_{n}-1\}$. Define $H_{s}$ to be the cyclic group on $O_{G,s}$. Define $G_{s}=\{g^{\frown}h:g\in G_{s-1},h\in H_{s}$}.
Stage $s=\langle n,m\rangle$ with $m>0$:
Check if $|W_{n,m}|>|W_{n,m-1}|$.
  • If no, then define $O_{G,s}=\{l_{s}\}$ and $G_{s}=\{g^{\frown}(l_{s}):g\in G_{s-1}\}$. Note, this gives that $H_{s}$ is the trivial group.
  • If yes, then take $N$ so that this is the $N$th time that $|W_{n,x}|>|W_{n,x-1}|$. That is, define
    $N=1+|\{x\in\mathbb{N}:0<x<m\,\&\,|W_{n,x}|>|W_{n,x-1}|\}|.$
    Define $O_{G,s}=\{l_{s},l_{s}+1,...,l_{s}+p_{n}^{N+1}-1\}$ and define $H_{s}$ to be the cyclic group on $O_{G,s}$. Define $t$ to be the stage that we had added an orbit of size $p_{n}^{N}$. We define
    $G_{s}=\{g_{i}^{\frown}h_{i}:0\leq i<p_{n}^{N}\,\&\,h_{i}\in H_{s}\,\&\,g_{i}\in G_{s-1}\text{ with }g_{i}(l_{t})-l_{t}\equiv h_{i}(l_{s})-l_{s}\mod p_{n}^{N}\}.$
For example, suppose we are at stage $1=\langle 0,1\rangle$. If $W_{0,1}=W_{0,0}=\emptyset$, then we will have
[…]
If $W_{0,1}\neq\emptyset$, then we will have
[…]
Verification
Since each $G_{s}$ is computable, it is clear that the tree $T_{G}$ is computable. Thus we need only show that $G\models\alpha_{n}$ if and only if $n\in\text{Fin}$.
Lemma 5
Let $s=\langle n,m\rangle$ with $|W_{n,m}|>|W_{n,m-1}|$. If $g\in G_{s}$ with $g^{p_{n}}=1_{G_{s}}$, then $g\upharpoonright{l_{s}}=1_{G_{s-1}}.$
Proof (of Lemma 5)
Let $N$ and $t$ be defined as they were at stage $s=\langle n,m\rangle$ of the construction. Note that $(g_{i}^{\frown}h_{i})^{p_{n}}=1_{G_{s}}$ if and only if $h_{i}^{p_{n}}=1_{H_{s}}$ and $g_{i}^{p_{n}}=1_{G_{s-1}}$. Since $h_{i}^{p_{n}}=1_{H_{s}}$, we must have that $h_{i}(l_{s})\equiv 0\mod p_{n}^{N}$. This gives that $g_{i}(l_{t})=l_{t}$, and so $g_{i}$ is the identity permutation when restricted to $O_{G,t}$. Similarly, we will get that $g_{i}$ is the identity permutation on $O_{G,r}$ for all $r$ of the form $r=\langle n,x\rangle$. In order for $g_{i}^{p_{n}}=1_{G_{s-1}}$, we must also have that $g_{i}$ is the identity permutation on $O_{G,r}$ for all $r$ that are not of the form $\langle n,x\rangle$ as all of these $O_{G,r}$ will either be of size 1, or of a size not divisible by $p_{n}$. Hence, we get that $g_{i}=g\upharpoonright l_{s}=1_{G_{s-1}}$. ∎
If $n\notin\text{Fin}$, then there will be infinitely many stages $s$ of the form $s=\langle n,m\rangle$ with $|W_{n,m}|>|W_{n,m-1}|$. Let $g\in G$. If $g^{p_{n}}=1$, then by Lemma 5 we have $g_{s-1}=g\upharpoonright l_{s-1}=1_{G_{s-1}}$ for each such $s$. However, if $g\in G$ was a witness to $\alpha_{n}$ then we would have by Lemma 2 that $G_{k}\models g_{k}\neq 1_{G_{k}}\,\&\,g_{k}^{p_{n}}=1_{G_{k}}$ (where $g_{k}=g\upharpoonright l_{k}$) for all but finitely many $k$, which is a contradiction. Hence, there is no witness to $\alpha_{n}$ in $G$ and so $G\models\neg\alpha_{n}$.
Now suppose that $n\in\text{Fin}$. We have that there is a least natural number $m$ such that $W_{n}$ gains no new elements after stage $m$. This gives that for $s=\langle n,m\rangle$, $|H_{s}|$ is a multiple of $p_{n}$, but no $H_{x}$ with $x>s$ will have $|H_{x}|$ divisible by $p_{n}$. Note that by our instructions, there will be an element $g\in G_{s}$ that is not the identity, but is of order $p_{n}$. Since we will never have $|W_{n,x}|>W_{n,x-1}$ for any $x>s$, we will have that there is an element of $G$ that is equal to $g$ on $G_{s}$, and is equal to the identity on all orbits higher than that of $G_{s}$. This element will be a witness to $\alpha_{n}$. ∎
So far we have only considered existential theories. We conclude by now expanding to entire first order theories for subgroups with orbit independence, which we show to be $\Delta^{0}_{2}$ relative to $\deg(T_{G})$ as a consequence of the following theorem of Feferman and Vaught.
Theorem 3.3 (Theorem 6.6 from [1])
Given any first order $\mathcal{L}$-sentence $\phi$, we can compute $n\in\mathbb{N}$ such that for every family $\{\mathcal{A}_{i}:i\in I\}$ of $\mathcal{L}$-structures there exists $J\subseteq I$ with $|J|\leq n$ such that if $\prod_{i\in I}\mathcal{A}_{i}\models\phi$, then $\prod_{i\in J^{\prime}}\mathcal{A}_{i}\models\phi$ for all $J^{\prime}$ with $J\subseteq J^{\prime}\subseteq I$. ∎
Corollary 2
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. Let $\alpha$ be any first order sentence in the language of groups. We have $G\models\alpha$ if and only if $G_{k}\models\alpha$ for all sufficiently large $k\in\mathbb{N}$.
Proof
Since $G$ has orbit independence, we have that $G\cong\prod_{i\in\mathbb{N}}H_{i}$. If $G\models\alpha$, then Corollary 3.3 gives that there is some finite $J\subset\mathbb{N}$ such that if $J^{\prime}\supseteq J$, then $\prod_{i\in J^{\prime}}H_{i}\models\alpha$. Thus, for all $k\geq\max(J)$ we have $G_{k}\models\alpha$. For the other direction note that if $G\nvDash\alpha$, then $G\models\neg\alpha$ and so the same reasoning gives that $G_{k}\models\neg\alpha$ for sufficiently large $k$. ∎
Theorem 3.4
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. The first order theory of $G$ is $\Delta^{0}_{2}$ relative to $\deg(T_{G})$.
Proof
Let $\text{Th}(G)$ denote the first order theory of $G$. By Corollary 2 we have that $\alpha\in\text{Th}(G)$ if and only if
$(\exists l)(\forall k>l)[G_{k}\models\alpha]$
which is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$. On the other hand, we have $\alpha\notin\text{Th}(G)$ if and only if
$(\exists l)(\forall k>l)[G_{k}\models\neg\alpha]$
which is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$. Hence, both $\text{Th}(G)$ and its complement are $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$ and so $\text{Th}(G)$ is $\Delta^{0}_{2}$ relative to $\deg(T_{G})$. ∎
This draws a strong distinction between the complexity of theories of profinite groups with and without orbit independence. Note that by the proof of Proposition 6 it is possible for just the existential theory of $G$ to be $\Sigma^{0}_{2}$ complete relative to $\deg(T_{G})$ when $G$ does not have orbit independence. However, the entire first order theory of $G$ will be $\Delta^{0}_{2}$ relative to $\deg(T_{G})$ when $G$ has orbit independence.